OpenStudy (anonymous):

if f is one to one find an equation for its inverse f(x)= 6/x-3 and f(x) = sqr(x+1) x > -1

myininaya (myininaya):

the first f(x)=6/(x-3) or 6/x-3?

OpenStudy (anonymous):

first is f(x) = 6/(x-3)

myininaya (myininaya):

ok so do you know how you would solve for x?

OpenStudy (anonymous):

solve for x, why do we need to solve for x, should we put (x-3)=0 ?

myininaya (myininaya):

no to find inverse all you have to do is solve for x and when you get done doing this replace x with f^{-1}(x) and replace f(x)=y with x ok y=6/(x-3) multiply both sides by x-3 y(x-3)=6 yx-3y=6 yx=6+3y x=(6+3y)/y f^{-1})(x)=(6+3x)/x

myininaya (myininaya):

2nd second one we have y=sqrt(x+1) square both sides y^2=x+1 subtract 1 y^2-1=x x=y^2-1 f^{-1}(x)=x^2-1

OpenStudy (anonymous):

thank you very much, really appreciate that